OpenStudy (anonymous):

The roots of the equation f(x)=0 are -1and -2. The roots f(-3x)=0 are ?

OpenStudy (anonymous):

The roots are going to be: \[x_1 = \frac{1}{3}, x_2 = \frac{2}{3}\] Because plugging those in f(-3x) yields: \[f(-3(\frac{1}{3})) = f(-1) = 0, f(-3(\frac{2}{3})) = f(-2) = 0\]